OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 15x + 9y = 240 6x + 3y = 90

OpenStudy (chaise):

15x + 9y = 240 6x + 3y = 90 Rearrange the equation - make one of the variables the subject. Seems easy to make y the subject in the second equation. 6x + 3y = 90 3y = 90 - 6x y = 30 - 2x SUBSTITUTE THIS VARIABLE back into the first example. 15x + 9y = 240 15x + 9(30-2x) = 240 15x + 270 - 18x = 240 -3x + 270 = 240 -3x = -30 x = 10 Substitute this answer back into any equation to solve for y. y = 30 - 2x y = 30- 2x10 y = 30-20 y = 10 x = 10 y = 10 (10,10) I think this was obvious from the word go, but there is your solution with working :) Hope I helped.

OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. -7x+7y=-7 and x+9y=61 can someone please help me with this one

OpenStudy (anonymous):

is anyone there to help me

OpenStudy (anonymous):

can someone please help me with solving using the elimination method for -3x+4y=1 and -6x+6y=-3. If the systime has no solution or an infinite number of solutions, state this.

OpenStudy (anonymous):

can someone help me